The charge which will flow  through a 200 ohm galvanometer connected to a 400 ohm circular coil of 1000 turns wound on a wooden stick 20 mm in diameter , if B = 0.012T parallal to the axis of stick decreased suddenly to zero
 
Ans 6.3 * 10^-6   C
 
Please give me a detailed solution i will ofcourse rate you


Grade:upto college level

1 Answers

ROSHAN MUJEEB
3 years ago
Here, total resistance of the circuit,R=200+400=600ΩR=200+400=600Ω
n=1000,D=20mm,r=10mm=10−2mn=1000,D=20mm,r=10mm=10-2m.
∴∴Area of cross-sectionA=πr2=π(10−2)2m2=10−4πm2A=πr2=π(10-2)2m2=10-4πm2
B1=0.0113T,B2=0,∴Db=B2−B1=−0.0113TB1=0.0113T,B2=0,∴Db=B2-B1=-0.0113T
Nowe=IR=dqdtRe=IR=dqdtR
Also,e=−ndϕdt=−nAdBdte=-ndϕdt=-nAdBdt
From (i) and (ii),dq=−nAdBR=−1000×10−4π×(−0.0113)600=+5.9×10−6C